OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = –1.7x + 8.5 17x + 10y = 85 –17x + 10y = 85 –17x + 10y = –85 17x + 10y = –85

OpenStudy (tyteen4a03):

Hint: -1.7 = -17/10. 8.5 = 85/10.